Hi can you help me on some simultaneous equations. I have found the answer to the variables but I need the working out Thanks.
3x-y=2 y=10
5x+2y=40 X=4
5x+3y=9 y=-2
7x+8y=5 x=3

Answer by checkley75(3666)

3X-Y=2, 3X-10=2, 3X=2+10, 3X=12, X=12/3, X=4 ANSWER.
5X+2Y=40, 5*4+2Y=40, 20+2Y=40, 2Y=40-20, 2Y=20, Y=20/2, Y=10 ANSWER.
I'LL LEAVE THE REST FOR YOU TO SOLVE.